# Optimal position-building strategies in Competition

### v1

@@ 4.3 Total Cost of trading
4.4 The balance of temporary and permanent impact and $\kappa$-regimes
@@ 5.1 Example calculation
In order to clarify the the exact nature of the trading costs, an example will be useful. Suppose $B$ is trading a trading a simple $\lambda$-scaled risk-neutral strategy $b(t)$:
$b(t)=\lambda\cdot t$
@@
$\text{Cost}(t)=\big{(}\dot{a}(t)+\dot{b}(t)\big{)}\cdot\dot{a}(t)+\kappa\cdot\big{(}a(t)+b(t)\big{)}\cdot\dot{a}(t)$
@@
Cumulative $\text{Cumulative cost}(t)=\int_{0}^{t}\big{(}\dot{a}(s)+\dot{b}(s)\big{)}\cdot\dot{a}(s)+\kappa\cdot\big{(}a(s)+b(s)\big{)}\cdot\dot{a}(s)\,\text{ds}$
@@
Figure 4: Example total costs for $a(t)$ in competition with $b(t)=t$ for $a(t)=t(t-c)/(1-c$. The left-hand column shows parabolic strategies $a(t)$ are denoted by solid lines and are all unit strategies with a variety of shapes from very risk-averse to very eager. The dotted lines in the right-hand column represent the risk-neutral strategy $b(t)$ and as we move down the rows its $\lambda$-scaling increases. Each plot in the right-hand column shows total costs of the strategies. One can see from the plots that the market impact caused by the $B$ makes risk-averse strategies the most costly as $B$ pushes the price up as time goes on. On the other hand when $A$ trades eagerly they can profit from over-buying the stock at lower prices and selling for a profit later. This is reflected, for example, in the most eager strategy’s cumulative cost peaking approximately sixty percent of the way through and then decreasing.
@@ 5.2 Motivating understanding the mix of temporary and permanent impact
The first question we ask is, how should trader $A$ go about choosing between t
@@
To get at an answer we start by observing that the primary concern for a trader regarding permanent impact is that as $B$ purchases the stock the price will remain higher throughout the buy-program and subsequent prices. In the absence of permanent impact, the trader’s only concern is minimizing temporary impact and the best way to do this in the case where $B$ is trading the risk-neutral strategy is to spread trades out as much as possible which clearly means trading the risk-neutral strategy as well (Figure 5 right plot)..
@@ 6.1 Almgren-Chriss optimal execution review
The setup In Almgren-Chriss is that at time $t=0$ a trader owns a certain quantity of and seeks to liquidate it entirely by time $t=T$. The liquidation trajectory is described by a twice-differentiable function $x(t)$ and the trader wishes to minimize a mix of temporary impact costs and risk as described by the following loss function [^3] The function could easily be called a cost function and be denoted $C$ but in keeping with the Euler-Lagrange equation we stick with $L$.:
$L=\dot{x}^{2}+\lambda\sigma^{2}x^{2}$ (5)
The strategy $x(t)$ that has minimal cost among all strategies is called the optimal strategy. To find the minimal strategy is a straightforward application of the Euler-Lagrange equation [^4] The Euler-Lagrange equation is ubiquitous, but see [GS+00] for a good overview.. Then the Euler-Lagrange equation is used to find the function $x(t)$ that minimizes a functional defined by means of the loss function $L(t,x(t),\dot{x}(t))$, where $\dot{x}(t)$ denotes the derivative of $x(t)$ with respect to $t$, the Euler-Lagrange equation is given by:
@@ 6.2 Permanent impact has no effect on optimal execution
$L=\dot{x}^{2}+\kappa\cdot x\cdot\dot{x}+\lambda\sigma^{2}x^{2}$ (9)
@@ 7 Non-equilibrium best-response strategies
To fashion this into a game, $A$ and $B$ simultaneously choose their strategies immediately prior to $t=0$ and agree to adhere perfectly to their respective strategies during the interval of time from $t=0$ to $t=1$. To begin our investigation we start by understanding how $B$ would trading knowing what strategy $B$ intends to trade. In the next few sections we derive non-equilibrium best-response strategies to particular strategies. The setup is as follows:
@@ 7.1 Best-response strategy to a risk-averse adversary
$L=\left(\dot{a}+\lambda\dot{b}\right)\cdot\dot{a}+\kappa\left(a+\lambda b\right)\cdot\dot{a}$ (11)
@@
$a(t)=-\frac{\lambda}{2}q(t)+\big{(}1+\frac{\lambda}{2}(q(1)-q(0))\big{)}\,t+\frac{\lambda}{2}q(0)$ (14)
Figure 8 shows an example of the best-response to a risk-averse position-building strategy. In each plot the red dotted strategy is $b(t)$, the risk-averse trader’s position-building strategy. In this situation the risk averse trader $B$ is simply trading without regard to the actions of trader $A$, while trader $A$ is trading optimally with respect to $B$. Notice that the blue lines are trading ahead of the risk-averse trader. The first column shows the situation where the risk-averse trader and the best-response trade the same amount. The second column shows where $\lambda=5$, that is, the risk-averse trader is acquiring five times as much stock as the best-response. In this case the best-response is to over-buy and then sell to the risk-averse trader as they continue to buy.
@@ 7.2 Best-response to a risk-neutral adversary
$a(t)=\left(1+-\frac{\lambda\kappa}{4}\right)t^{2}+\frac{\lambda\kappa}{4}t$
@@ 7.3 Best-response to an eager adversary
[…]
@@
$a(t)=\frac{e^{-\lambda t}\left(e^{\lambda}(\lambda-\kappa)-te^{\lambda t}(\lambda-\kappa+2)+e^{\lambda(t+1)}(\kappa-\lambda+t(\lambda-\kappa+2))\right)}{2\left(e^{\lambda}-1\right)}$
@@
Examining the plots we see something curious. The best-response to the eager trader counter-intuitively begins by selling short and then later re-buys the stock. Since the eager trader is rapidly pushing the price of the stock up, can this possibly be correct? The answer comes down to precisely precisely how the model deals with market impact.
@@
$1,0.1$ 2.00, $,\,2.00$ 2.00 0.11, $,\,0.11$ 0.11 2.09, 2.112.092.112.09,\,2.112.09 , 2.11
$1,0.3$ 2.00, $,\,2.00$ 2.00 0.28, 0.270.280.270.28,\,0.270.28 , 0.27 2.26, 2.272.262.272.26,\,2.272.26 , 2.27
$1,2.5$ 2.11, 2.002.112.002.11,\,2.002.11 , 2.00 2.55, 2.702.552.702.55,\,2.702.55 , 2.70 4.66, 4.704.664.704.66,\,4.704.66 , 4.70
$1,10$ 4.03, 2.004.032.004.03,\,2.004.03 , 2.00 7.13, 10.87.1310.87.13,\,10.87.13 , 10.8 11.2, 12.811.212.811.2,\,12.811.2 , 12.8
$1,25$ 14.8, 2.0014.82.0014.8,\,2.0014.8 , 2.00 2.46, 27.12.4627.12.46,\,27.12.46 , 27.1 17.2, 29.117.229.117.2,\,29.117.2 , 29.1
$3,0.1$ 2.52, 4.002.524.002.52,\,4.002.52 , 4.00 0.36, 0.270.360.270.36,\,0.270.36 , 0.27 2.88, 4.272.884.272.88,\,4.272.88 , 4.27
$3,0.3$ 2.53, 4.002.534.002.53,\,4.002.53 , 4.00 0.89, 0.660.890.660.89,\,0.660.89 , 0.66 3.42, 4.663.424.663.42,\,4.663.42 , 4.66
$3,2.5$ 3.55, 4.003.554.003.55,\,4.003.55 , 4.00 7.05, 6.647.056.647.05,\,6.647.05 , 6.64 10.6, $,\,10.6$ 10.6
$3,10$ 19.0, 4.0019.04.0019.0,\,4.0019.0 , 4.00 3.57, 26.63.5726.63.57,\,26.63.57 , 26.6 22.5, 30.622.530.622.5,\,30.622.5 , 30.6
$3,25$ 105, 4.001054.00105,\,4.00105 , 4.00 114.3, 66.4114.366.4-114.3,\,66.4- 114.3 , 66.4 9.09, 70.49.0970.4-9.09,\,70.4- 9.09 , 70.4
$10,0.1$ 89.0, 11.089.011.0-89.0,\,11.0- 89.0 , 11.0 2.93, 0.952.930.952.93,\,0.952.93 , 0.95 86.1, 12.086.112.0-86.1,\,12.0- 86.1 , 12.0
$10,0.3$ 89.0, 11.089.011.0-89.0,\,11.0- 89.0 , 11.0 7.3, 2.47.32.47.3,\,2.47.3 , 2.4 81.7, 13.481.713.4-81.7,\,13.4- 81.7 , 13.4
$10,2.5$ 82.8, 11.082.811.0-82.8,\,11.0- 82.8 , 11.0 61.3, 23.861.323.861.3,\,23.861.3 , 23.8 21.5, 34.821.534.8-21.5,\,34.8- 21.5 , 34.8
$10,10$ 11.0, $,\,11.0$ 11.0 95.0, $,\,95.0$ 95.0 106.0, $,\,106.0$ 106.0
$10,25$ 536, 11.053611.0536,\,11.0536 , 11.0 513, 238513238-513,\,238- 513 , 238 24, 2492424924,\,24924 , 249
$25,0.1$ 1771, 26.0177126.0-1771,\,26.0- 1771 , 26.0 16.8, 2.4516.82.4516.8,\,2.4516.8 , 2.45 1754, 28.5175428.5-1754,\,28.5- 1754 , 28.5
$25,0.3$ 1771, 26.0177126.0-1771,\,26.0- 1771 , 26.0 41.7, 6.141.76.141.7,\,6.141.7 , 6.1 1729, 32.1172932.1-1729,\,32.1- 1729 , 32.1
$25,2.5$ 1753, 26.0175326.0-1753,\,26.0- 1753 , 26.0 385, 61.338561.3385,\,61.3385 , 61.3 1368, 87.3136887.3-1368,\,87.3- 1368 , 87.3
$25,10$ 1483, 26.0148326.0-1483,\,26.0- 1483 , 26.0 1108, 24511082451108,\,2451108 , 245 376, 271376271-376,\,271- 376 , 271
$25,25$ 26.0, $,\,26.0$ 26.0 613, $,\,613$ 613 639, $,\,639$ 639
Table 1: The temporary and permanent impact costs of the best-response (first number) and risk-neutral (second number) strategies in competition with an eager trader as depicted in Figure 12.
@@ 7.4 Temporary impact revisited
Compare this, however, to the situation in which a trader $A$ is trading $a(t)$ in competition with another trader, say $B$, trading $\lambda b(t)$. In this situation, the market impact cost to $A$ will depend on the net trading in the stock. It will be a profit (a negative cost) whenever $\text{sgn}(\dot{a}(t)+\lambda{\dot{b}(t)})\neq\text{sgn}(\dot{a}(t))$. That is, whenever the trading direction of $A$ and $B$ in aggregate is the opposite sign of $A$’s then $A$’s temporary impact will be a profit for $A$. We illustrate this in a small table:
@@ 8.1 Joint best-response strategies in general
Using the same concepts as in sections 7.1 to 7.3 we derive equations defining best-response strategies for $A$ and $B$ position-building in competition where $B$ trades a $\lambda$-scaled strategy. First we derive $A$’s best-response strategy for $B$ noting the loss function is as follows:
$L_{a}$ $=(\dot{a}+\lambda\dot{b})\cdot\dot{a}+\kappa(a+\lambda b)\cdot\dot{a}$ (17)
$L_{b}$ $=(\dot{a}+\lambda\dot{b})\cdot\lambda\dot{b}+\kappa(a+\lambda b)\cdot\lambda\dot{b}$ (18)
Using the Euler-Lagrange equation Eq. 6 for $L_{a}$ and $L_{b}$ respectively yield the equilibrium equations with boundary conditions as given:
@@ 8.2 Extremal kappa regimes
Before studying equilibrium in general we examine two extreme cases in order to build intuition. Fix $\lambda>$ and. as usual, consider two strategies trading in competition: a unit strategy $A$ and a $\lambda$-scaled strategy $B$. In equilibrium, the strategies satisfy the equations Eqs. 19 to 20 above.
@@
$\ddot{b}$ $=-\frac{1}{2\lambda}$ (22)
@@
$L_{a}$ $=\kappa(a+\lambda b)\cdot\dot{a}$
$L_{b}$ $=\kappa(a+\lambda b)\cdot\lambda\dot{b}$
@@
and with the position-building boundary conditions there is no solution. One way to understand what is going on is the look at how the unit trader strategies look at what happens to the solutions $a,b$ for the equilibrium equations Eqs. 19 to 20 for increasingly large values of $\kappa$, to understanding the limiting behavior of the equations as $\kappa\to\infty$. Here are plots for this scenario:
Figure 13: Solutions to the equilibrium equations Eqs. 19 to 20 as $\kappa$ grows large for a unit trading $A$ and a $\lambda$-scaled trader $B$ trading five times as much as $A$. One can see that the unit trader $a(t)$’s strategy grows toward a strategy of instantly acquiring the three-times its target position instantly and then selling down to target an instant before completion time, while $b(t)$ tends toward acquiring approximately sixty-percent immediately and the rest an instant before completion time.

### v2

@@ 4.3 Total Cost of trading
4.4 The balance of temporary and permanent impact: market imppact regimes
@@ 5.1 Example calculation
In order to clarify the the exact nature of the trading costs, an example will be useful. Suppose $B$ is trading the $\lambda$-scaled risk-neutral strategy $b(t)$:
$b_{\lambda}(t)=\lambda\cdot t$
@@
$\text{Cost}(t)=\big{(}\dot{a}(t)+\dot{b}_{\lambda}(t)\big{)}\cdot\dot{a}(t)+\kappa\cdot\big{(}a(t)+b_{\lambda}(t)\big{)}\cdot\dot{a}(t)$
@@
Cumulative $\text{Cumulative cost}(t)=\int_{0}^{t}\big{(}\dot{a}(s)+\dot{b}_{\lambda}(s)\big{)}\cdot\dot{a}(s)+\kappa\cdot\big{(}a(s)+b_{\lambda}(s)\big{)}\cdot\dot{a}(s)\,\text{ds}$
@@
Figure 4: Example total costs for $a(t)$ in competition with $b_{\lambda}(t)=t$ for $a(t)=\frac{1}{1-c}(2t-c)$. The left-hand column shows parabolic strategies $a(t)$ are denoted by solid lines and are all unit strategies with a variety of shapes from very risk-averse to very eager. The dotted lines in the right-hand column represent the risk-neutral strategy $b_{\lambda}(t)$ and as we move down the rows its $\lambda$-scaling increases. Each plot in the right-hand column shows total costs of the strategies. One can see from the plots that the market impact caused by the $B$ makes risk-averse strategies the most costly as $B$ pushes the price up as time goes on. On the other hand when $A$ trades eagerly they can profit from over-buying the stock at lower prices and selling for a profit later. This is reflected, for example, in the most eager strategy’s cumulative cost peaking approximately sixty percent of the way through and then decreasing.
@@ 5.2 Motivating understanding the mix of temporary and permanent impact
The first question we ask is, how should trader $A$ go about choosing between various strategies.
@@
To get at an answer we start by observing that the primary concern for a trader regarding permanent impact is that as $B$ purchases the stock the price will remain higher throughout the buy-program and subsequent prices. In the absence of permanent impact, the trader’s only concern is minimizing temporary impact and the best way to do this in the case where $B$ is trading the risk-neutral strategy is to spread trades out as much as possible which clearly means trading the risk-neutral strategy as well (Figure 5 right plot).
@@ 6.1 Almgren-Chriss optimal execution review
The setup In Almgren-Chriss is that at time $t=0$ a trader owns a certain quantity of and seeks to liquidate it entirely by time $t=T$. The liquidation strategy’s trajectory is described by a twice-differentiable function $x(t)$; the trader wishes to minimize some combination of temporary impact cost and risk as described by the following loss function [^3] The function could easily be called a cost function and be denoted $C$ but in keeping with the Euler-Lagrange equation we stick with $L$.:
$L(t;x)=\dot{x}^{2}+\lambda\sigma^{2}x^{2}$ (5)
This is a function of time, $t$, and we note its dependence on the strategy $x$ as well. The strategy $x(t)$ that has minimal cost among all strategies is called the optimal strategy. To find the minimal strategy is a straightforward application of the Euler-Lagrange equation [^4] The Euler-Lagrange equation is ubiquitous, but see [GS+00] for a good overview.. Then the Euler-Lagrange equation is used to find the function $x(t)$ that minimizes a functional defined by means of the loss function $L(t,x(t),\dot{x}(t))$, where $\dot{x}(t)$ denotes the derivative of $x(t)$ with respect to $t$, the Euler-Lagrange equation is given by:
@@ 6.2 Permanent impact has no effect on optimal execution
$L(x)=\dot{x}^{2}+\kappa\cdot x\cdot\dot{x}+\lambda\sigma^{2}x^{2}$ (9)
@@ 7 Non-equilibrium best-response strategies
To fashion this into a game, $A$ and $B$ simultaneously choose their strategies immediately prior to $t=0$ and agree to adhere perfectly to their respective strategies during the interval of time from $t=0$ to $t=1$. To begin our investigation we start by understanding how $B$ should trade knowing what strategy $A$ will trade. In the next few sections we derive non-equilibrium best-response strategies to particular strategies. The setup is as follows:
@@ 7.1 Best-response strategy to a risk-averse adversary
$L(a,b)=\left(\dot{a}+\lambda\dot{b}\right)\cdot\dot{a}+\kappa\left(a+\lambda b\right)\cdot\dot{a}$ (11)
@@
$a(t)=-\dfrac{\lambda}{2}q(t)+\big{(}1+\dfrac{\lambda}{2}(q(1)-q(0))\big{)}\,t+\dfrac{\lambda}{2}q(0)$ (14)
Summary: We can summarise the best-response to a risk-averse trader more neatly as follows. The best response to the $\lambda$-scaled risk-averse strategy $b(t;{\lambda,\sigma})=\frac{\sinh(\sigma t)}{\sinh(t)})$ is given by a strategy $a(t;{\kappa,\!\lambda,\!\sigma\!})$, where we explicitly signal the dependence on the risk-aversion parameter $\sigma$, scaling $\lambda$ and market impact coefficient $\kappa$:
$a(t;{\kappa,\!\lambda,\!\sigma\!})=\dfrac{\lambda}{2}(q_{\kappa,\sigma\!}(0)-q_{\kappa,\sigma\!}(t))+\big{(}1+\dfrac{\lambda}{2}(q_{\kappa,\sigma\!}(1)-q_{\kappa,\sigma\!}(0))\big{)}\,t$ (15)
where $q_{\sigma,\lambda}(t)$ is an auxiliary function depending on $\sigma$ and $\kappa$:
$q_{{\kappa,\sigma\!}}(t)=\frac{\sinh(\sigma\,t)}{\sinh(\sigma)}+\frac{\kappa}{\sigma}\cdot\frac{\cosh(\sigma\,t)}{\sinh(\sigma)}$ (16)
Plots of the best-response to a risk-averse trader: Figure 8 shows an example of the best-response to a risk-averse position-building strategy. In each plot the solid red line represents the strategy $b(t)$, the risk-averse trader’s position-building strategy. In this situation the risk averse trader $B$ is simply trading without regard to the actions of trader $A$, while trader $A$ is trading optimally with respect to $B$. Notice that the blue lines are trading ahead of the risk-averse trader. The first row shows the situation where the risk-averse trader and the best-response target the same quantity (i.e., $\lambda=1$). The second column shows where $\lambda=5$, that is, the risk-averse trader is acquiring five times as much stock as the best-response. In this case the best-response is to over-buy and then sell to the risk-averse trader as they continue to buy.
@@ 7.2 Best-response to a risk-neutral adversary
$a(t)=\left(1+\frac{\lambda\kappa}{4}\right)t-\frac{\lambda\kappa}{4}t^{2}$
@@ 7.3 Best-response to an eager adversary
$b(t;{\lambda,\sigma})=\frac{e^{-\sigma t}-1}{e^{-\sigma}-1},\qquad\text{eager $\sigma$-scaled, $\sigma$ eagerness}$
@@
$b(t;{\kappa,\!\lambda,\!\sigma\!})=\frac{e^{\sigma(-t)}\left(\lambda e^{\sigma}(\sigma-\kappa)-te^{\sigma t}((\lambda+2)\sigma-\kappa\lambda)+e^{\sigma+\sigma t}(-\lambda\sigma+\kappa(\lambda-\lambda t)+(\lambda+2)\sigma t)\right)}{2\left(e^{\sigma}-1\right)\sigma}$ (19)
@@
Examining the plots we see something curious. The best-response to the eager trader counter-intuitively begins by selling short and then later re-buys the stock. Since the eager trader is rapidly pushing the price of the stock up, can this possibly be correct? The answer comes down to how the model precisely deals with market impact.
Table 1: The temporary and permanent impact costs of the best-response (first number) and risk-neutral (second number) strategies in competition with an eager strategy b(t) with eagerness $\sigma=4$, as depicted in Figure 12.
@@
$,0.1$ 0.10 ${2.12},2.0$ ${0.12},0.11$ ${2.24},2.11$
$1,0.25$ ${2.10},2.0$ ${0.30},0.27$ ${2.41},2.27$
$1,2.50$ ${1.99},2.0$ ${2.84},2.70$ ${4.82},4.70$
$1,10$ ${2.81},2.0$ ${8.72},10.82$ ${11.53},12.82$
$1,25$ ${10.14},2.0$ ${8.64},27.05$ ${18.78},29.05$
$3,0.10$ ${2.63},4.0$ ${0.39},0.27$ ${3.02},4.27$
$3,0.25$ ${2.60},4.0$ ${0.96},0.66$ ${3.56},4.66$
$3,2.50$ ${2.93},4.0$ ${7.82},6.64$ ${10.75},10.64$
$3,10$ ${15.08},4.0$ ${7.76},26.57$ ${22.84},30.57$
$3,25$ ${90.40},4.0$ ${-98.1},66.4$ ${-7.8},70.4$
$10,0.10$ ${-56.6},11$ ${1.88},0.95$ ${-54},12$
$10,0.25$ ${-55.0},11$ ${4.61},2.38$ ${-50.33},13$
$10,2.50$ ${-21.1},11$ ${32.71},23.75$ ${11.66},35$
$10,10$ ${214.7},11$ ${-48.29},95.0$ ${166},106$
$10,25$ ${1253},11$ ${-1016},238$ ${237},249$
$25,0.10$ ${-580},26$ ${5.51},2.5$ ${-574},28$
$25,0.25$ ${-563},26$ ${13.5},6.1$ ${-549},32$
$25,2.50$ ${-245},26$ ${91},61$ ${-154},87$
$25,10$ ${1582},26$ ${-226},245$ ${1356},271$
$25,25$ ${8776},26$ ${-3510},613$ ${5267},639$
@@ 7.4 Temporary impact revisited
Compare this, however, to the situation in which a trader $A$ is trading $a(t)$ in competition with another trader, say $B$, trading $\lambda b(t)$. In this situation, the market impact cost to $A$ will depend on the net trading in the stock. It will be a profit (a negative cost) whenever $\text{sgn}(\dot{a}(t)+\lambda{\dot{b}(t)})\neq\text{sgn}(\dot{a}(t))$. That is, whenever the trading direction of $A$ and $B$ is in aggregate the opposite sign of $A$’s then $B$’s temporary impact will create a profit opportunity for $A$. We illustrate this in a small table:
@@ 8.1 Joint best-response strategies in general
Using the same concepts as in sections 7.1 to 7.3 we derive equations defining best-response strategies for $A$ and $B$ position-building in competition where $B$ trades a $\lambda$-scaled strategy. First we note that the loss functions for $A$ and $B$ are given as:
$L(a,b)$ $=(\dot{a}+\lambda\dot{b})\cdot\dot{a}+\kappa(a+\lambda b)\cdot\dot{a}$ (20)
$L(b,a)$ $=(\dot{a}+\lambda\dot{b})\cdot\lambda\dot{b}+\kappa(a+\lambda b)\cdot\lambda\dot{b}$ (21)
Now applying the Euler-Lagrange equation Eq. 6 for $L(a,b)$ and $L(b,a)$ simultaneously yields the equilibrium equations:
@@ 8.2 Extremal kappa regimes
Before studying equilibrium in general we examine two extreme cases in order to build intuition. Fix $\lambda>0$ and. as usual, consider two strategies trading in competition: a unit strategy $A$ and a $\lambda$-scaled strategy $B$. In equilibrium, the strategies satisfy the equations Eqs. 22 to 23 above.
@@
$\ddot{b}$ $=-\frac{1}{2\lambda}\ddot{a}$ (25)
@@
$L(a,b)$ $=\kappa(a+\lambda b)\cdot\dot{a}$
$L(b,a)$ $=\kappa(a+\lambda b)\cdot\lambda\dot{b}$
@@
and with the position-building boundary conditions there is no solution. One way to understand what is going on is the look at how the unit trader strategies look at what happens to the solutions $a,b$ for the equilibrium equations Eqs. 22 to 23 for increasingly large values of $\kappa$, to understanding the limiting behavior of the equations as $\kappa\to\infty$. Here are plots for this scenario:
Figure 13: Solutions to the equilibrium equations Eqs. 22 to 23 as $\kappa$ grows large for a unit trading $A$ and a $\lambda$-scaled trader $B$ trading five times as much as $A$. One can see that the unit trader $a(t)$’s strategy grows toward a strategy of instantly acquiring the three-times its target position instantly and then selling down to target an instant before completion time, while $b(t)$ tends toward acquiring approximately sixty-percent immediately and the rest an instant before completion time.